HUBBLE SPACE TELESCOPE DAILY REPORT #2597 PERIOD COVERED: 0000Z (UTC) 04/06/00 - 0000Z (UTC) 04/07/00 Daily Status Report as of 098/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 8214 (The Evolution of Globular Cluster Systems In Early-Type Galaxies) The WF/PC-2 was used to measure structural parameters {core radii and concentrations} of globular clusters in three early-type galaxies using deep, four-point dithered observations. The observations completed nominally. 1.2 Completed Two Sets of WF/PC-2 8460 (Cycle 8 Supplemental Darks pt2/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations completed with no reported problems. 1.3 Completed WF/PC-2 8203 (Dwarf Busting in Abell 868) The WF/PC-2 was used to to identify and characterize the morphological nature of the dwarf galaxy population{s} responsible for the steep upturn seen in Abel 868's luminosity function. While similar upturns have been reported in many clusters, the specifics of the dwarf population remain unknown as these objects cannot be resolved from the ground. What type of dwarf galaxies are they ? Is there more than one population contributing ? How are they clustered ? These observations are intended to answer these questions. The observations were completed, and no problems were reported. 1.4 Completed Two Sets of WF/PC-2 8457 (UV Earth Flats) The WF/PC-2 was used to obtain a series of Ultraviolet Earth flats. The observations were completed with no reported problems. 1.5 Completed Two Sets of STIS/CCD 8410 (STIS CCD Hot Pixel Annealing C8) The Space Telescope Imaging Spectrograph (CCD) was used to assess the effectiveness of the CCD hot pixel annealing process by measuring the dark current behavior before and after annealing and by searching for any window contamination effects. In addition CTE performance is examined by looking for traps in a low signal level flat. The observations were completed as planned, and no anomalies were reported. 1.6 Completed WF/PC-2 8444 (WF/PC-2 CYCLE-8 Internal Monitor) The WF/PC-2 was used to make a bias, and an internal flat, observation as part of a Cycle-8 monitoring program to monitor the health of the cameras. A variety of internal exposures are obtained in order to provide a monitor of the integrity of the CCD camera electronics in both bays {gain 7 and gain 15}, a test for quantum efficiency in the CCDs, and a monitor for possible buildup of contaminants on the CCD windows. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed Two Sets of WF/PC-2 8230 (Snapshot Survey of Dynamically Close Galaxy Pairs from z=0.1 to z=0.5) The WF/PC-2 was used to make snapshot survey observations of galactic cluster Gal-Clus-0. This is to obtain snapshot images of a large, well-defined, and dynamically selected sample of close galaxy pairs at moderate redshift. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed Two Sets of STIS/CCD 8549 (POMS Test Proposal: Non-Scripted Parallel Proposal Continuation I) The Space Telescope Imaging Spectrograph (CCD) was used to make low galactic latitude, non-scripted parallel observations as part of a POMS test proposal. The observations were completed as planned, and no anomalies were reported. 1.9 Completed Two Sets of WF/PC-2 8448 (Cycle 8 INTFLAT Sweeps, VISFLAT Mini-Sweep, and Linearity Test) The WFPC2 was used to monitor the pixel to pixel flatfield response and the VISFLAT lamp degradation, as well as detect any possible changes due to contamination. As described in 2.1 and HSTAR 7568, the acquisition for the final iteration of this proposal failed, and the take data flag remained down, affecting three observations. Otherwise, the observations completed nominally. 1.10 Completed STIS/CCD/MA1/MA2 7764 (Activity on the Edge of Convection: The Atmosphere of Canopus {F0 Ib-II}) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was to be used to make observations of the star Canopus (a.k.a., HD45348 or Alpha Car), the nearest {96 pc} supergiant star. Canopus is an early F supergiant with an extremely thin convection zone, yet it shows significant X-ray emission from a very hot corona. A wavelength calibration image was also to be taken. As described in 2.1 and HSTAR 7568, the acquisition for this proposal failed, and the take data flag remained down. Twenty-one observations were affected. 1.11 Completed WF/PC-2 8090 (POMS Test Proposal: WF-II Parallel Archive Proposal Continuation) The WF/PC-2 was to be used in the parallel mode to make some low galactic latitude archive observations. As described in 2.1 and HSTAR 7568, the acquisition for this proposal failed, and the take data flag remained down. Five observations were affected. 1.12 Completed WF/PC-2 8193 (Comparing the Hosts of High-Z Radio-Quiet Quasars to Lyman Break Galaxies) The WF/PC-2 was used to try to determine the underlying causes of the rapid cosmological evolution of the quasar population that is a mystery around which is woven many different strands of contemporary cosmology. The proposal completed nominally. 1.13 Completed STIS/CCD 8437 (Dark Monitor-Part 2 C8) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks. The proposal completed nominally. 1.14 Completed STIS/CCD 8439 (Bias Monitor-Part 2 C8) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4, to build up high-S/N superbiases and track the evolution of hot columns. The proposal completed with no reported problems. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 7 Successful: 6 Scheduled Re-acquisitions: 8 Successful: 7 The acquisition scheduled at 097/145753Z failed to RGA mode at 097/150151Z when the search radius limit was exceeded on both FGS-1 and FGS-2. FHST maps taken before the failure were nominal. A map performed at 097/1537Z showed small vehicle errors. The subsequent re-acquisition at 097/1636Z also failed. HSTAR 7568 was written. The proposals described in 1.9, 1.10, and 1.11 were affected. Per HSTAR 7570, the acquisition at 097/235030 defaulted to fine lock backup on FGS-3 only when the scan step limit was exceeded on FGS-2. Two proposals, WF/PC-2 8114 and STIS 8439, both occurring during the period of the next report, may have been affected. 2.2 FHST Updates: Scheduled: 18 Successful: 18 2.3 Operations Notes: The load for the Target of Opportunity Observation, STIS proposal 8747, was uplinked mid-morning yesterday. Using ROP SR-1A, the SSR EDAC error counter was cleared twice. STIS error flags were cleared at 097/1629Z per ROP NS-12. This relates to the acquisition problems with the proposal described in 1.10.. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /CAW